use keplers 3rd law to determine the mass of the sun. that is the equation we derived in class which states that the orbital period squared is proportional to the orbital radius cubed. you can tell if you got the right answer by looking it up on wikipedia-sear for sun.

Keplers third law: (T/2T) (r3/GM)

Where T is the period and r is the distance between the satellite and the much larger mas.

Distance from the Earth to the sun = 1.5*10^11 m

Earth year = 365.25 days
Answer #1

we have Kepler third Law

T2/(4\pi2)=r3/(GM )

Now

M=(4\pi2r3)/(GT2)

we have T=365.25 days

T=365.25\times24\times60\times60 =31557600 s

r=1.5\times1011 m , G=6.67\times10-11

Now putting all values

M=(4\times(3.14)2\times(1.5\times1011 )3) / ((6.67\times10-11 )\times(31557600 )2)

M=2.0 \times1030 Kg